OpenStudy (anonymous):

Cosxtanx/cscx = 1/4 Solve for x

OpenStudy (anonymous):

\[\frac{ cosxtanx }{ cscx } = \frac{ 1 }{ 4 }\]\[\frac{ sinx }{ cscx } = \frac{ 1 }{ 4 }\]\[\sin^2x = \frac{ 1 }{ 4 }\]\[\sqrt{\sin^2x} = \sqrt{\frac{ 1 }{ 4 }}\]\[sinx = \pm 1/2\]\[x = 30°, 150°, 210°, 330°\]